OpenStudy (anonymous):

9z^2=-6z+15 solve the equation and check your solution

OpenStudy (anonymous):

9z² = -6z + 15 9z² + 6z - 15 = 0 9z² - 9z + 15z - 15 = 0 9z(z - 1) + 15(z - 1) = 0 (9z + 15)(z-1) = 0 (z-1) --> z = 1 (9z + 15) --> z = -15/9 --> z = -5/3 So z = 1 or z = -5/3